# Lattice simulations of axion-U(1) inflation: gravitational waves, magnetic fields, and black holes

### v1

@@ 6 Curvature perturbations and its probability distribution function
In the previous analytical study [35, 36, 37], the formation of the PBHs from the large spatial fluctuations of the axion field generated by the backreaction of the gauge field on the axion dynamics has been studied. To estimate the abundance of PBHs, the authors used the Press-Schechter approach [100, 101, 102]. For this, one needs to know the probability distribution function (PDF) of the curvature perturbation variable, $\zeta=-H\delta\phi/\dot{\bar{\phi}}$. The PDF of $\zeta$ will be the same as the PDF of spatial fluctuations, $\delta\phi$, of the axion field. In these studies, this PDF has been assumed to be a $\chi^{2}$-distribution due to the quadratic nature of the backreaction term of the gauge field in the scalar field dynamics. Using this, the authors in [35, 36] examined the constraints on $\alpha/f$ by using the constraints on PBHs abundances from BBN and CMB and obtained an upper bound on $\alpha/f$. The obtained values of the upper bound are $\alpha/f\leq 115$ from [35] and $\alpha/f\leq 130$ from [36]. The obtained upper bound on $\alpha/f$ in these studies is such that this setup of axion-U(1) inflation is in the strong backreaction regime. Here we test this assumption and study the PDF of the axion field fluctuations in our simulations.
@@
To study the evolution of the PDF of the axion field’s spatial fluctuations, we use 3D snapshots from our simulations. At the beginning of our simulations, these fluctuations follow a Gaussian random field as set by the initial conditions. We normalize by the root mean square of the fluctuations in the axion field, $\sigma_{\phi}$, and plot the PDF of $\delta\phi/\sigma_{\phi}\equiv(\phi-\bar{\phi})/\sigma_{\phi}$. Initially, it follows a standard normal distribution. In figure 6, we present the evolution of the PDFs for three different coupling strengths, $\alpha/f=35$ (top panel), $\alpha/f=60$ (mid panel) and $\alpha/f=75$ (bottom panel). As evident from the top panel of figure 6, the PDF of $\delta\phi/\sigma_{\phi}$ is initially Gaussian and evolves into a $\chi^{2}$-distribution with three degrees of freedom. However, in the case of strong coupling strength, the PDF starts as a Gaussian, transitions to a $\chi^{2}$ distribution at an intermediate time, and reverts to a distribution closer to Gaussian at later times.
@@
Considering a distribution that is less non-Gaussian than the $\chi^{2}$-distribution will result in a smaller PBH abundance [^8] To accurately estimate PBH abundances, it is necessary to know the PDF of the smoothed curvature perturbations [35, 36]. Since the PDF of the smoothed curvature perturbations will be similar to, or even less non-Gaussian than, that of the unsmoothed perturbations, the constraints derived from the smoothed perturbations will be comparable to or less stringent than those from the unsmoothed perturbations.. This will relax the constraints on $\alpha/f$. We will present a detailed study of the PBHs constraints in a future study.
@@ 7 Discussion and conclusion
The large-scale gauge fields produced in this model offer a possible explanation for intergalactic magnetic fields. We have explored this for small to large coupling strengths, $\alpha/f$, and find that the resulting magnetic fields are above the lower bound on intergalactic magnetic fields inferred from blazar observations [67] for $\alpha/f\geq 60$. For the coupling strength $\alpha/f=75$, the resulting magnetic field strength is $1.2\times 10^{-14}\,{\rm G}$ with a coherence length of $0.19\,{\rm pc}$. The lower bound on the magnetic field strength from blazar observations for this coherence length is $7.2\times 10^{-15}\,{\rm G}$, based on the relation $B_{\rm bound}=10^{-17}\,{\rm G}\left(L/0.1\,{\rm Mpc}\right)^{1/2}$, as discussed in section 5.2.
Furthermore, we examine the PDF of the produced axion fluctuations. Understanding this PDF is crucial for studying the formation of PBHs in this model. Previous analytical studies considered it to be a $\chi^{2}$-distribution with one or two degrees of freedom. In our study, we initialize axion fluctuations as a Gaussian random field. We find that the spatial fluctuations indeed transition to a $\chi^{2}$-distribution at later times, but with three degrees of freedom. However, as the backreaction of the gauge field becomes significant, the system starts to resemble a Gaussian distribution. Incorporating this information in studies constraining the axion-gauge coupling from PBH abundance bounds will relax the existing constraints on this coupling strength.
@@ Acknowledgments
We would like to thank Angelo Caravano, Yutong He, Oksana Iarygina, and Alberto Roper Pol for the discussions on this project. We are also indebted to Kai Schmitz and Oleksandr Sobol for providing the datasets used to compare with the results from the gradient expansion formalism for the homogeneous axion case as well as for the discussions on this comparison. R.S. and A.V. were supported by the Czech Science Foundation (GAČR), project 24-13079S. A.B. acknowledges funding from the Swedish Research Council (Vetenskapsrådet) under grant No. 2019-04234 and the National Science Foundation under Grant No. AST-2307698, and the NASA ATP Award 80NSSC22K0825. We also thank the Swedish National Allocations Committee for providing computing resources at the Center for Parallel Computers at the Royal Institute of Technology in Stockholm and the National Supercomputer Centre (NSC) in Linköping. Some simulations in this work were performed on the “Phoebe” computer cluster at CEICO/FZU. R.S. is grateful to Josef Dvořáček for assistance with setting up the Pencil Code on Phoebe. K.S. was partially supported by the Alexander Von Humboldt Foundation through the Carl Friedrich von Siemens Research Award. He thanks Volker Springel his host, as well as Torsten Ensslin and Eichiro Komatsu for warm hospitality during his stay at the Max-Planck Institute for Astrophysics, Garching.
@@
The source code used for the numerical solutions of this study, the Pencil Code, along with the module special/backreact_infl used in the present study, are freely available at https://github.com/pencil-code/pencil-code/. The numerical data and input files are available on http://norlx65.nordita.org/~brandenb/projects/axion-U1/.
@@ Appendix A Implementation in the Pencil Code
As discussed in section 2, we use the Pencil Code for our simulations. To simulate the axion-U(1) inflation, we use the special modules backreact_inf.f90 and disp_current.f90. The $\phi$ and FLRW background equations are solved in the ‘backreact_infl.f90’ module, while the $\mathbf{E}$ and $\mathbf{A}$ equations are solved in the ‘disp_current.f90’ module. By default, we use the Weyl gauge, but the Lorenz gauge can also be used by setting the llongitudinalE logical to false and the llorenz_gauge_disp logical to true. The governing equations in both Lorenz and Weyl gauges are described in the section A.1 and section A.2, respectively. In addition to the $\mathbf{E}$ and $\mathbf{A}$ variables, we also solve for $A_{0}$ in the Lorenz gauge. Additionally, we also solve for the production of gravitational waves using the equations in the gravitational module gravitational_waves_hTXk.f90; see ref. [110] for details. The Pencil Code employs a sixth-order finite difference method for spatial derivatives and a third-order Runge-Kutta method for time stepping.
@@ Appendix C Comparison with the GEF results
Importantly, the difference in gauge field energy density between GEF and our simulation remains small after the time of backreaction even in the case in which we initialize at $N_{i}=-3.14$ for $\alpha/f=90$. Therefore, while choosing better initial conditions could improve consistency with GEF, it would not significantly impact the predictions for magnetic fields and gravitational waves.
Appendix D Spectrum of scalar field fluctuations
At the end of section 3.3, we compared with the results for runs D and E. Here, we provide the spectra of the scalar field fluctuation $P_{\delta\phi}$. The left- and right-hand panels of figure 9 show those spectra for runs D and E, respectively. The light to dark shades represent different times, with each curve separated by $\Delta N=0.5$. They agree well with those of ref. [53].
Figure 9: The spectrum of scalar field fluctuations is shown, with the left panel representing run D and the right panel representing run E. The curves progress in time from lighter to darker shades, each separated by $\Delta N=0.5$.

### v2

@@ 6 Curvature perturbations and its probability distribution function
In the previous analytical study [41, 42, 43], the formation of the PBHs from the large spatial fluctuations of the axion field generated by the backreaction of the gauge field on the axion dynamics has been studied. To estimate the abundance of PBHs, the authors used the Press-Schechter approach [133, 134, 135]. For this, one needs to know the probability distribution function (PDF) of the curvature perturbation variable, $\zeta=-H\delta\phi/\dot{\bar{\phi}}$  [^12] This expression for $\zeta$ is valid only within the linear approximation.. The PDF of $\zeta$ will be the same as the PDF of spatial fluctuations, $\delta\phi$, of the axion field. In these studies, this PDF has been assumed to be a $\chi^{2}$-distribution due to the quadratic nature of the backreaction term of the gauge field in the scalar field dynamics. Using this, the authors in [41, 42] examined the constraints on $\alpha/f$ by using the constraints on PBHs abundances from BBN and CMB and obtained an upper bound on $\alpha/f$. The obtained values of the upper bound are $\alpha/f\leq 115$ from [41] and $\alpha/f\leq 130$ from [42]. The obtained upper bound on $\alpha/f$ in these studies is such that this setup of axion-U(1) inflation is in the strong backreaction regime. A recent numerical study [76] found that the non-Gaussianity of curvature perturbations is strongly suppressed, and the PDF of $\zeta$ is nearly Gaussian. Here we also test the assumption made in previous analytical studies [41, 42] by analyzing the PDF of the axion field fluctuations in our simulations.
@@
To study the evolution of the PDF of the axion field’s spatial fluctuations, we use 3D snapshots from our simulations. At the beginning of our simulations, these fluctuations follow a Gaussian random field as set by the initial conditions. We normalize by the root mean square of the fluctuations in the axion field, $\sigma_{\phi}$, and plot the PDF of $\delta\phi/\sigma_{\phi}\equiv(\phi-\bar{\phi})/\sigma_{\phi}$. Initially, it follows a standard normal distribution. In figure 7, we present the evolution of the PDFs for three different coupling strengths, $\alpha/f=35$ (top panel), $\alpha/f=60$ (mid panel) and $\alpha/f=75$ (bottom panel). In this figure, the dashed yellow line in the left column represents a standard normal distribution, while the dashed magenta line in the right column corresponds to a $\chi^{2}$-square distribution with three degrees of freedom. The variable $\delta\phi/\sigma_{\phi}$ has zero mean and unit variance. To compare it with a $\chi^{2}$-square distribution with different degrees of freedom, we use a transformed variable $\sqrt{2d}\,\delta\phi/\sigma_{\phi}+d$ with a mean of $d$ and a variance of $2d$. This transformation is used because a $\chi^{2}$-square distributed variable with $d$ degrees of freedom has a mean $d$ and a variance of $2d$. In figure 7, we show the $\chi^{2}$-distribution for $d=3$, as it provides the best fit. As is evident from the top panel of figure 7, the PDF of $\delta\phi/\sigma_{\phi}$ is initially Gaussian and evolves into a $\chi^{2}$-distribution with three degrees of freedom. However, in the case of strong coupling strength, the PDF starts as a Gaussian, transitions to a $\chi^{2}$ distribution at an intermediate time, and reverts to a distribution closer to Gaussian at later times. It is to be noted that the PDF at $N=0.2$ for the case $\alpha/f=75$ has less power in the tails compared to the $\chi^{2}$-distribution. However, it is still not close to a Gaussian distribution. The exact reason for this behavior is not clear, but it is expected that the PDF at later times, $N>0.2$, will have a more Gaussian-like distribution.
@@
Considering a distribution that is less non-Gaussian than the $\chi^{2}$-distribution will result in a smaller PBH abundance [^13] To accurately estimate PBH abundances, it is necessary to know the PDF of the smoothed curvature perturbations [41, 42]. We expect that the PDF of the smoothed curvature perturbations will be similar to or even less non-Gaussian than that of the unsmoothed perturbations. This is because smoothing involves averaging fluctuations using a window function that assigns different weights to points at different separations. Hence, a non-Gaussian distribution will gradually approach a Gaussian one due to the central limit theorem. Taking this into consideration, the constraints derived from the smoothed perturbations will be comparable to or less stringent than those from the unsmoothed perturbations. . This will relax the constraints on $\alpha/f$. We will present a detailed study of the PBHs constraints in a future study.
@@ 7 Discussion and conclusion
The large-scale gauge fields produced in this model offer a possible explanation for intergalactic magnetic fields. We have explored this for small to large coupling strengths, $\alpha/f$, and find that the resulting magnetic fields are only marginally above the lower bound on intergalactic magnetic fields inferred from blazar observations [94] for $\alpha/f\geq 60$. For the coupling strength $\alpha/f=90$, the resulting magnetic field strength is $1.9\times 10^{-14}\,{\rm G}$ with a coherence length of $0.25\,{\rm pc}$. The lower bound on the magnetic field strength from blazar observations for this coherence length is $1.6\times 10^{-14}\,{\rm G}$, based on the relation $B_{\rm bound}=1.8\times 10^{-17}\,{\rm G}\left(L/0.2\,{\rm Mpc}\right)^{1/2}$, as discussed in section 5.2. We note that the generated magnetic field strengths could be reduced upon incorporating the effects of charged currents produced through interactions between gauge fields and fermions, e.g., via the Schwinger effect. We leave a detailed investigation of this effect for future work.
Furthermore, we examine the PDF of the produced axion fluctuations. Understanding this PDF is crucial for studying the formation of PBHs in this model. Previous analytical studies considered it to be a $\chi^{2}$-distribution with one or two degrees of freedom. In our study, we initialize axion fluctuations as a Gaussian random field. We find that the spatial fluctuations indeed transition to a $\chi^{2}$-distribution at later times, but with three degrees of freedom. However, as the backreaction of the gauge field becomes significant, even if only after inflation, the system starts to resemble a Gaussian distribution as also previously demonstrated in ref. [76] for the case $\alpha/f=125$, when the backreaction is important already during inflation. Incorporating this information in studies constraining the axion-gauge coupling from PBH abundance bounds will relax the existing constraints on this coupling strength.
@@ Acknowledgments
We would like to thank Angelo Caravano, Valerie Domcke, Yutong He, Oksana Iarygina, and Alberto Roper Pol for the discussions on this project. We are also indebted to Kai Schmitz and Oleksandr Sobol for providing the datasets used to compare with the results from the gradient expansion formalism for the homogeneous axion case as well as for the discussions on this comparison. R.S. and A.V. were supported by the Czech Science Foundation (GAČR), project 24-13079S. A.B. acknowledges funding from the Swedish Research Council (Vetenskapsrådet) under grant No. 2019-04234 and the National Science Foundation under Grant No. AST-2307698, and the NASA ATP Award 80NSSC22K0825. We also thank the Swedish National Allocations Committee for providing computing resources at the Center for Parallel Computers at the Royal Institute of Technology in Stockholm and the National Supercomputer Centre (NSC) in Linköping. Some simulations in this work were performed on the “Phoebe” computer cluster at CEICO/FZU. R.S. is grateful to Josef Dvořáček for assistance with setting up the Pencil Code on Phoebe. K.S. was partially supported by the Alexander Von Humboldt Foundation through the Carl Friedrich von Siemens Research Award. He thanks Volker Springel his host, as well as Torsten Ensslin and Eichiro Komatsu for warm hospitality during his stay at the Max-Planck Institute for Astrophysics, Garching.
@@
The source code used for the numerical solutions of this study, the Pencil Code, along with the module special/backreact_infl used in the present study, are freely available at https://github.com/pencil-code/pencil-code/. The numerical data and input files are available on Zenodo; see ref. [142].
@@ Appendix A Implementation in the Pencil Code
As discussed in section 2, we use the Pencil Code for our simulations. To simulate the axion-U(1) inflation, we use the special modules backreact_inf.f90 and disp_current.f90. The $\phi$ and FLRW background equations are solved in the ‘backreact_infl.f90’ module, while the $\mathbf{E}$ and $\mathbf{A}$ equations are solved in the ‘disp_current.f90’ module. By default, we use the Weyl gauge, but the Lorenz gauge can also be used by setting the llongitudinalE logical to false and the llorenz_gauge_disp logical to true. The governing equations in both Lorenz and Weyl gauges are described in the section A.1 and section A.2, respectively. In addition to the $\mathbf{E}$ and $\mathbf{A}$ variables, we also solve for $A_{0}$ in the Lorenz gauge. Additionally, we also solve for the production of GWs using the equations in the gravitational module gravitational_waves_hTXk.f90; see ref. [143] for details. The Pencil Code employs a sixth-order finite difference method for spatial derivatives and a third-order Runge-Kutta method for time stepping.
@@ Appendix C Comparison with the GEF results
Importantly, the difference in gauge field energy density between GEF and our simulation remains small after the time of backreaction even in the case in which we initialize at $N_{i}=-3.14$ for $\alpha/f=90$. Therefore, while choosing better initial conditions could improve consistency with GEF, it would not significantly impact the predictions for magnetic fields and GWs.
Appendix D Spectrum of scalar field fluctuations and of helicity of the magnetic component of the gauge field
At the end of section 3.3, we compared with the results for runs D and E. Here, we provide the spectra of the scalar field fluctuation $P_{\delta\phi}$ and the spectra of the helicity of the magnetic component of the gauge field, $H_{M}(k)$ defined as $\int{\rm d}k~{}H_{M}(k)=\langle(\mathbf{A}\cdot\mathbf{B})\rangle$. The left- and right-hand panels of figure 10 show those spectra for runs D and E, respectively. The light to dark shades represent different times, with each curve separated by $\Delta N=0.5$. They agree well with those of ref. [77].
Figure 10: The spectrum of scalar field fluctuations and of the helicity of the magnetic component of the gauge field is shown, with the left panel representing run D and the right panel representing run E. The curves progress in time from lighter to darker shades, each separated by $\Delta N=0.5$ for the upper panel and $\Delta N=0.1$ for the lower panel.